The minimum deviations suffered by red, yellow and violet beams passing through an equilateral transparent. Prism are 38.4°, 38.7° and 39.2° respectively. Calculate the dispersive power of the medium.

Grade:11

2 Answers

Aditi Chauhan
9 years ago
Given that, δ base r = 38.4°, δ base y = 38.7° and δ base v = 39.2° Dispersive power = μ base v - μ base r/ μ base y – 1 = (μ base v - 1) – (μ base v - 1) – (μ base r - 1)/( μ base y - 1) = (δ base v/A) - (δ base r/A)/ (δ base v/A) [∵ δ = (μ – 1) A] = δ base v - δ base r/ δ base y = 39.2 – 38.4/38.7 = 0.0204
Navjyot Kalra
9 years ago
Sol. Given that, δ base r = 38.4°, δ base y = 38.7° and δ base v = 39.2° Dispersive power = μ base v - μ base r/ μ base y – 1 = (μ base v - 1) – (μ base v - 1) – (μ base r - 1)/( μ base y - 1) = (δ base v/A) - (δ base r/A)/ (δ base v/A) [∵ δ = (μ – 1) A] = δ base v - δ base r/ δ base y = 39.2 – 38.4/38.7 = 0.0204